OpenStudy (anonymous):

Refresh my memory. a negative plus a negative=? a negative subtract a negative=? a negative plus a positive=? a negative times a negative=? a positive times a positive=? a negative divided by a negative=? a negative subtract a positive=? a positive times a negative=?

Directrix (directrix):

Suggestion: Make up a problem to illustrate each of the 8 situations you posted. That way, it will be easier to help, I think.

OpenStudy (anonymous):

-8+-8=? -8-(-8)=? -8+8=? 8+8= -8-8=? -8*8=? -8*(-8)=? 8*8=? -8/8=? -8/-8=? 8/8=?

OpenStudy (anonymous):

-8+-8= -16

OpenStudy (anonymous):

-8-(-8=0

OpenStudy (anonymous):

8+8=16

OpenStudy (anonymous):

hope that helped

OpenStudy (anonymous):

yes thank you